OpenStudy (anonymous):

Let f(x) = 3x +2 and g(x) = 7x + 6. Find f • g and its domain.

OpenStudy (anonymous):

A. 6x2 + 4x + 42; all real numbers except x = - 2/3 B. 6x2 + 4x + 42; all real numbers C. 21x2 + 32x + 12; all real numbers D. 21x2 + 32x + 12; all real numbers except x = -6/7

OpenStudy (sirm3d):

the dot is multiplication, am i right?

OpenStudy (anonymous):

yes C?

OpenStudy (anonymous):

it is multiplication, unless it is actually a small o then i'm blind lol. anyhow... it is C that's right

OpenStudy (anonymous):

:D